# CPEG: Leveraging Consistency Policy with Consensus Guidance for Multi-agent Exploration

### v1

@@ Abstract
Efficient exploration is crucial in cooperative multi-agent reinforcement learning (MARL), especially in sparse-reward settings. However, due to the reliance on the unimodal policy, existing methods are prone to falling into the local optima, hindering the effective exploration of better policies. Furthermore, tackling multi-agent tasks in complex environments requires cooperation during exploration, posing substantial challenges for MARL methods. To address these issues, we propose a Consistency Policy with consEnsus Guidance (CPEG), with two primary components: (a) introducing a multimodal policy to enhance exploration capabilities, and (b) sharing the consensus among agents to foster agent cooperation. For component (a), CPEG incorporates a Consistency model as the policy, leveraging its multimodal nature and stochastic characteristics to facilitate exploration. Regarding component (b), CPEG introduces a Consensus Learner to deduce the consensus on the global state from local observations. This consensus then serves as a guidance for the Consistency Policy, promoting cooperation among agents. The proposed method is evaluated in multi-agent particle environments (MPE) and multi-agent MuJoCo (MAMuJoCo), and empirical results indicate that CPEG not only achieves improvements in sparse-reward settings but also matches the performance of baselines in dense-reward environments.
@@ I Introduction
Recent years have witnessed a growing body of applications in cooperative multi-agent reinforcement learning (MARL), such as multi-robot tasks [1] and autonomous driving [2]. Despite these successful applications, cooperative MARL still faces challenges in exploration due to limitations in the policy class regarding multi-modality and the necessity for cooperation in multi-agent systems (MAS) [3].
In single-agent RL, the exploration quality heavily depends on the chosen policy class of agents [4, 5, 6], with inappropriate policy classes potentially leading to local optima. This issue becomes more pronounced in MARL owing to the complex interaction between agents. In MARL, popular methods [7, 8, 9] commonly formulate the continuous policy as a unimodal density function, typically a Gaussian distribution. While computationally efficient, these policies can significantly weaken the exploration, as the sampled actions tend to be concentrated around the modality. Besides, the unimodal policy is prone to converging towards a suboptimal policy due to the lack of expressiveness, overfitting the behavior of other agents. To address these challenges, some methods [10, 4, 11] explore alternative policy classes to enhance exploration. Although these methods improve exploration to some extent, they often exhibit limitations in practice. For example, Gaussian mixture models can only cover a limited number of modes, and normalizing flow methods, while able to compute density values, suffer from numerical instability due to their determinant dependence. With the growing prevalence of diffusion models [12], a series of works [13, 14, 15] apply them as a powerful and multi-modal policy class, primarily in the context of single-agent offline RL. However, the diffusion model is time-intensive, involving multiple sampling steps (e.g., 1000 steps), which makes the training and execution computationally expensive, especially for the online MARL heavily depends on sampling from environments. In order to accelerate the speed of the sampling process, a novel model, the consistency model is designed to map any point at any time step back to the start of the trajectory, based on the probability flow ordinary differential equation (PF-ODE) [12].
@@
Figure 1: An example of cooperative exploration. The two-agent arm requires collaborative exploration to reach four targets at different locations. In a sparse reward setting, agents must reach all targets before receiving any reward, making exploration more challenging.
Besides, the cooperation among agents during exploration is critical in complex environments [16, 17, 3]. To exemplify, consider the Reacher4 task depicted in Fig. 1, wherein two agents are required to coordinate their swinging motions to ensure that the end-effector of the manipulator contacts one of the four targets. In a sparse reward setting, agents must reach all targets before receiving any reward, making exploration more challenging. In this scenario, the agents need to coordinate their actions to set the entire robotic arm in motion, instead of engaging in independent and meaningless actions. Additionally, the agents need to adapt to the multimodality brought by the four targets, avoiding premature convergence to a unimodal policy and achieving simultaneous exploration of multiple targets. This challenge can be tackled through the use of communication-based MARL techniques [18]. However, these techniques introduce challenges such as selecting appropriate information to transmit and additional bandwidth requirements. Recently, some studies [19, 20] leverage a shared consensus among agents to promote cooperation. At each timestep, although the local observations of each agent are unique, they represent different aspects of the same global state. In the given scenario, observations of the agents only cover their own information such as angular velocity, which serves as partial representations of the global state. However, these individual observations are merely projections of the state. Consequently, our objective is to extract the global state information through consensus, thereby fostering cooperation.
@@
Inspired by the aforementioned observations, we introduce a novel framework tailored to MARL, termed the Consistency Policy with consEnsus Guidance (CPEG), aimed at facilitating efficient cooperative exploration. In this work, we adopt a powerful and effective model, the consistency model as the policy class, with the objective of harnessing its stochastic nature and expressiveness to explore in a multimodal manner. Compared with diffusion models, consistency models enable efficient one-step generation while retaining the advantages of multi-step iterative sampling. For cooperative exploration, we introduce a discrete consensus representation as the guidance for the consistency policy. Specifically, we employ a codebook from VQ-VAE [21] for a discrete, distinguishable consensus representation. This allows agents to derive the same estimation of the global state from a unified consensus codebook, thereby guiding cooperative behaviors. To strike a balance between exploration and exploitation during training, some studies [6, 22] use the mask to decide on exploration or exploitation. Building on the insights from these work, we design a guidance mask to intermittently drop the guidance with probability. Additionally, during the initial phase of training, generative policies may produce unreliable actions, which can affect the training of the policy. To address this, we introduce a self-reference mechanism during the training phase, leveraging past successful experiences to constrain the actions generated by the consistency policy, thus facilitating exploration.
@@
We empirically evaluate the proposed method on two distinct environments: the multiple-particle environment (MPE) [7] and multi-agent MuJoCo (MAMuJoCo) [23]. Across all experiments, we explore both dense-reward settings and sparse-reward settings, where the latter poses significant challenges for exploration as agents receive rewards solely upon completion of a specified task. The results demonstrate that CPEG exhibits notable superiority over competitive baselines in sparse-reward settings and achieves comparable performance in dense-reward settings, thereby underscoring the effectiveness of our approach.
@@
    We introduce a Consistency Policy for agents, capable of completing the diffusion process in a single step, to explore the environment in a multimodal way. To the best of our knowledge, our study represents the first effort in leveraging consistency policies in MARL.
@@
    We propose a Consensus Learner to infer the global consensus from local observations, aimed at guiding efficient cooperative exploration by sharing the same consensus between agents.
@@
    We evaluate our method on MPE and MAMuJoCo with dense and sparse reward settings. The results indicate that CPEG performs comparably to state-of-the-art (SOTA) algorithms in dense reward environments, while in sparse reward environments, which are more challenging for exploration, our algorithm outperforms SOTA algorithms by 20%.
@@ II-A Exploration in MARL
As one of the most critical issues in reinforcement learning, the exploration has drawn great attention in single-agent RL research. This can be even more challenging in complex environments with sparse reward [3]. In the multi-agent domain, several works [24, 16, 17] have emerged to address exploration challenges. MAVEN [24], for instance, aims to maximize the mutual information between the trajectories and latent variables, thereby enabling the learning of diverse exploration policies. EITI and EDTI [16], on the other hand, are designed to capture the influence of one agent’s behaviors on others, encouraging agents to visit states that influence the behavior of others. CMAE [17] adopts a shared exploration goal derived from multiple projected state spaces. Beyond these considerations, recent works highlight that the choice of policy class employed by the agents can significantly impact exploration dynamics [4, 5]. However, these efforts primarily concentrate on single-agent domains. In this paper, we highlight the equally crucial role of policy classes in decision-making within multi-agent systems. In our proposed approach, CPEG, we introduce an innovative policy class, the consensus-guided consistency policy, specifically tailored to enhance exploration in MARL.
@@ III-B Consistency Models
The diffusion model [12] addresses the multimodal distribution matching problem using a stochastic differential equation (SDE), while the consistency model [36] tackles a comparable probability flow ordinary differential equation (PF-ODE): ${\mathrm{d}x_{\tau}}/{\mathrm{d}\tau}=-\tau\nabla\log p_{\tau}(x)$, where $\tau\in[0,T]$, $T>0$ is a fixed constant. Here, $p_{\tau}(x)=p_{\text{data}}(x)\otimes\mathcal{N}(0,\tau^{2}I)$ denotes the distribution of data $x_{\tau}$ at step $\tau$, where $\otimes$ denotes the convolution operation and $p_{\text{data}}(x)=p_{0}(x)$ represents the raw data distribution. The reverse process occurs along the solution trajectory $\{\hat{x}_{\tau}\}_{\tau\in[\epsilon,T]}$, with $\epsilon$ being a small constant close to 00, employed for handling the numerical issues at the boundary. To accelerate the sampling process of a diffusion model, the consistency model reduces the required number of sampling steps to a significantly smaller value without substantially compromising model generation performance Specifically, it approximates a parameterized consistency function $f_{\theta}:(x_{\tau},\tau)\to x_{\epsilon}$, which is defined as a map from the noisy sample $x_{\tau}$ at step $\tau$ back to the original sample $x_{\epsilon}$, in contrast to the step-by-step denoising function $p_{\theta}(x_{\tau-1}|x_{\tau})$ as the reverse diffusion process in diffusion model. The consistency function possesses the property of self-consistency, meaning its outputs are consistent for arbitrary pairs of $(x_{\tau},\tau)$ that belong to the same PF-ODE trajectory. Besides, for modeling the conditional distribution with condition variable $c$, the consistency function is modified to $f_{\theta}(c,x_{\tau},\tau)$, representing a slight deviation from the original consistency model.
@@ IV Method
Figure 2: The overall framework of the proposed CPEG. Consensus-guided diffusion policies facilitate cooperative exploration among multiple agents. The consistency policy utilizes the agent’s observation and Gaussian noise as inputs, generating actions under the guidance of masked consensus. The consensus learner consists of an observation encoder, consensus codebook, and state decoder, which learns discrete consensus representations by reconstructing states. During the training phase, in addition to policy learning, the self-reference mechanism provides gradient back-propagation based on the disparity between the agent’s action and those from the reference buffer, thereby imposing a policy constraint.
In this section, we introduce CPEG, a novel approach that integrates the consistency policy guided by consensus between agents for exploration. As illustrated in Fig. 2, our framework consists of three principal components: a) a consistency policy that generates actions from Gaussian noise, b) a discrete consensus representation with masks, which serves as a guidance for cooperation, and c) a self-reference mechanism to constrain consistency policies. Detailed discussions on each component are provided in the subsequent sections.
@@ IV-A Consistency Policy
Notation
Given the presence of two types of timesteps in this work, one for the diffusion process and one for reinforcement learning, we use superscripts $\tau\in[0,T]$ to denote diffusion timesteps and subscripts $t\in\{1,\dots,T_{e}\}$ to denote environment timesteps.
@@
where $\mu(\cdot,\cdot)$ representing the drift coefficient, $\sigma(\cdot)$ representing the diffusion coefficient, and $\{w^{\tau}\}_{\tau\in[0,T]}$ denoting the standard Brownian Motion. Beginning with $u_{a}^{T}$, the diffusion policy aims to recover the original action $u_{a}^{0}=u_{a}$ by solving a reverse process from $T$ to 00 using the reverse-time SDE:
@@
Although this diffusion policy can be used for exploration, it is time-intensive and makes the policy inference unacceptably slow, especially in tasks that contain multiple agents. To address this issue, we introduce a consistency policy [36, 34] based on Eq. 3:
@@
where the initial action $u_{a}^{\tau_{n}}\sim\mathcal{N}(0,{\tau_{n}}I)$, and $\theta_{a}$ represents the parameters of the policy $\pi$ of the agent $a$. We discretize the diffusion horizon $[\epsilon,T]$ into a predetermined sequence $\{\tau_{n}|n\in N\}$ of length $N$ for determining the solution trajectory of action. Moreover, $\tau_{n}$ is the $n$-th sub-intervals of the time horizon, with $n\sim\mathcal{U}(1,N-1)$. The consistency function $F_{\theta_{a}}(o_{a},u_{a}^{\tau},\tau)$ is a trainable deep neural network that accepts the observation $o_{a}$ of agent $a$ as an input and generates an action with dimensions matching $u_{a}^{\tau}$. $c_{skip}$ and $c_{out}$ are differentiable functions, with $c_{\mathrm{skip}}(\epsilon)=1$, $c_{\mathrm{out}}(\epsilon)=0$, ensuring the consistency process remains differentiable at a fixed, small positive timestep $\epsilon$.
@@
Policy Learning
To train the consistency policy, we adopt an off-policy optimization approach similar to MADDPG [7]. Following the approach of Wang et al. [13], we integrate the Q-value function into the consistency model, training it to preferentially sample actions with higher values:
@@
Figure 3: The framework of the Consensus Learner. It operates differently during the training and execution process.
IV-B Consensus Guidance
To guide cooperation in multi-agent exploration, we introduce a consensus learner, where agents sharing the same consensus demonstrate cooperative behaviors. We first formulate a set of $K$ discrete consensuses $\Psi\coloneqq\{\psi_{1},\dots,\psi_{K}\}$, where $K\in\mathbb{N}$ is a tunable hyperparameter. Each consensus $\psi_{k}$ is defined by a tuple $\left\langle e_{\psi_{k}},I_{\psi_{k}}\right\rangle$, where $k\in\{1,\dots,K\}$ is the identity of consensus and $e_{\psi_{k}}\in\mathbb{R}^{m}$ is the code (or latent embedding) of consensus $\psi_{k}$. $I_{\psi_{k}}$ is the set of agents who have the same consensus $\psi_{k}$, and each agent can only have one consensus at each timestep: $I_{\psi_{i}}\cap I_{\psi_{j}}=\varnothing,\cup_{j}I_{\psi_{j}}=I$ for $i,j\in\{1,2,\dots,K\}$ and $i\neq j$. To represent consensus, we design a consensus leaner with VQ-VAE [21], as shown in Fig. 3. The motivations for this include: a) The discrete codebook in VQ-VAE can effectively extract state features with similar semantics, providing better separation than traditional continuous representation methods. b) The use of discrete feature representation helps mitigate the impact of unavoidable noise on exploration in stochastic environments. c) The discretization process ensures a distinct association, where each agent is linked to a single, specific consensus.
The consensus learner comprises four primary components: an observation encoder, a consensus codebook, a state decoder, and a consensus mask. We assume the consensus can be deduced from the current global state $s$. However, for distributed execution in CTDE, each agent $a$ can only access its local observation $o_{a}$. Considering $o_{a}=\Omega(s,a)$ as a projection of the global state $s$, we design a two-phase mechanism to deduce consensus from local observations. To deduce the consensus from local observations, the consensus learner has two modes during the training and execution phases: during the training phase, the consensus learner processes observations from all agents, obtains discrete consensus representations for each agent’s observation, and uses them to reconstruct the global state; in the execution phase, it processes agent $a$’s observation to obtain the consensus for guiding policy generation of action $u_{a}$. In this manner, the consensus learner develops the capacity to infer consensus based on local observations.

### v2

# CPIG: Leveraging Consistency Policy with Intention Guidance for Multi-agent Exploration

@@ Abstract
Efficient exploration is crucial in cooperative multi-agent reinforcement learning (MARL), especially in sparse-reward settings. However, due to the reliance on the unimodal policy, existing methods are prone to falling into the local optima, hindering the effective exploration of better policies. Furthermore, in sparse-reward settings, each agent tends to receive a scarce reward, which poses significant challenges to inter-agent cooperation. This not only increases the difficulty of policy learning but also degrades the overall performance of multi-agent tasks. To address these issues, we propose a Consistency Policy with Intention Guidance (CPIG), with two primary components: (a) introducing a multimodal policy to enhance the agent’s exploration capability, and (b) sharing the intention among agents to foster agent cooperation. For component (a), CPIG incorporates a Consistency model as the policy, leveraging its multimodal nature and stochastic characteristics to facilitate exploration. Regarding component (b), we introduce an Intention Learner to deduce the intention on the global state from each agent’s local observation. This intention then serves as a guidance for the Consistency Policy, promoting cooperation among agents. The proposed method is evaluated in multi-agent particle environments (MPE) and multi-agent MuJoCo (MAMuJoCo). Empirical results demonstrate that our method not only achieves comparable performance to various baselines in dense-reward environments but also significantly enhances performance in sparse-reward settings, outperforming state-of-the-art (SOTA) algorithms by 20%.
@@ I Introduction
Recent years have witnessed a growing body of applications in cooperative multi-agent reinforcement learning (MARL), such as multi-robot tasks [1] and autonomous driving [2]. Despite these successful applications, cooperative MARL still faces challenges in exploration due to limitations in the policy class regarding multi-modality and the necessity for cooperation in multi-agent systems (MAS) [3, 4].
@@
Figure 1: An example of cooperative exploration. The two-agent arm requires collaborative exploration to reach four targets at different locations. In a sparse reward setting, agents should reach all targets before receiving any reward, making exploration and cooperation more challenging. The top-right corner illustrates a visualization of the multimodal joint policy.
In single-agent RL, the exploration quality heavily depends on the chosen policy class of agents [5, 6, 7], with inappropriate policy classes potentially leading to local optima. This issue becomes more pronounced in MARL owing to the complex interaction between agents. In MARL, popular methods [8, 9, 10, 11] commonly formulate the continuous policy as a unimodal density function, typically a Gaussian distribution. While computationally efficient, these policies can significantly weaken the exploration, as the sampled actions tend to be concentrated around one modality. Besides, the unimodal policy is prone to converging towards a suboptimal policy due to the lack of expressiveness, overfitting the behavior of other agents. To address these challenges, some methods [12, 5, 13] explore alternative policy classes to enhance exploration. Although these methods improve exploration to some extent, they often exhibit limitations in practice. For example, Gaussian mixture models can only cover a limited number of modes, and normalizing flow methods [14], while able to compute density values, suffer from numerical instability due to their determinant dependence. With the growing prevalence of diffusion models [15], a series of works [16, 17, 18] apply them as a powerful and multi-modal policy class, primarily in the context of single-agent offline RL. However, the diffusion model is time-intensive, involving multiple diffusion steps (e.g., 1000 steps), which makes the training and execution computationally expensive, especially for the online MARL heavily depends on sampling from environments. In order to accelerate the sampling process, we employ a novel generative model, the consistency model, which is designed to map any diffusion point at any time step back to the start of the generative process based on the probability flow ordinary differential equation (PF-ODE) [15]. In multi-agent systems, the cooperation among agents during exploration is critical in complex environments [19, 20, 4]. To exemplify, consider the Reacher4 task depicted in Fig. 1, wherein two agents are required to coordinate their swinging motions to ensure that the end-effector of the manipulator contacts one of the four targets. In this scenario, the agents need to coordinate their actions to set the entire robotic arm in motion, instead of engaging in independent and meaningless actions. In a sparse reward setting, agents must reach all targets before receiving any reward, making exploration more challenging. Additionally, as illustrated in the top-right corner of the example, we visualize how this environment requires policies to handle a multimodal distribution. The agents must adapt to the multimodality introduced by the four targets, avoiding premature convergence to a unimodal policy and enabling simultaneous exploration of multiple targets. This challenge can be tackled through the use of communication-based MARL techniques [21]. However, these techniques introduce challenges such as selecting appropriate information to transmit and additional bandwidth requirements.
Recently, some studies [22, 23, 24] leverage a shared feature among agents, such as intention or consensus, to promote cooperation. This concept stems from the fact that at each timestep, while the local observations of individual agents are unique, they collectively represent different facets of the same global state. In the given scenario, though observations of the agents only cover their local information such as angular velocity, this information also serves as a partial representation of the global state. Therefore, these individual observations can be viewed as projections of the global state. In this paper, one of our primary objectives is to extract the state information through an intention learner, thereby fostering cooperation.
@@
Building on the aforementioned observations, we propose a novel framework for MARL, termed Consistency Policy with Intention Guidance (CPIG), designed to facilitate efficient cooperative exploration. In this framework, we adopt the consistency model as the policy class, leveraging its powerful expressiveness and stochastic nature to facilitate multimodal exploration. Compared to diffusion models, consistency models enable efficient one-step generation while preserving the benefits of multi-step iterative sampling. To enhance cooperative exploration, we incorporate discrete intention as guidance for the consistency policy. Specifically, we employ a codebook for a discrete, distinguishable intention representation. This allows agents to derive the same estimation of the global state from a unified intention codebook, thereby guiding cooperative behaviors. Besides, to strike a balance between exploration and exploitation during training, some studies [7, 25] use the mask to decide on exploration or exploitation. Building on insights from these prior works, we propose a guidance mask that periodically drops the guidance with a certain probability. When guidance is dropped, agents tend to engage in individual random exploration. In contrast, with guidance, their behavior shifts toward a focus on exploitation and cooperative behaviors. Additionally, during the initial phase of training, generative policies may produce unreliable actions, which can affect the training of the policy. To address this, we introduce a self-reference mechanism during the training phase, leveraging past successful experiences to constrain the actions generated by the policy, thus facilitating exploration.
@@
We conduct empirical evaluations of the proposed method on two distinct environments: the multiple-particle environment (MPE) [8] and multi-agent MuJoCo (MAMuJoCo) [26]. Across all experiments, we explore both dense-reward settings and sparse-reward settings. The sparse-reward settings present significant challenges for exploration, as agents receive rewards only upon completing a specified task. The results demonstrate that CPIG exhibits notable superiority over competitive baselines in sparse-reward settings and achieves comparable performance in dense-reward settings, thereby underscoring the effectiveness of our approach.
@@
    We propose a Consistency Policy for agents, which can complete the diffusion process in a single step, enabling them to explore the environment in a multimodal manner. To the best of our knowledge, our method is the first to leverage consistency policies in MARL.
@@
    We propose an Intention Learner to infer the global intention from local observations, aimed at guiding efficient cooperative exploration by sharing the same intention between agents.
@@
    We evaluate our method on MPE and MAMuJoCo with dense and sparse reward settings. The results indicate that CPIG performs comparably to baselines in dense reward environments, while in sparse reward environments, which are more challenging for exploration, our algorithm outperforms SOTA algorithms by 20%.
@@ II-A Exploration in MARL
As one of the most critical issues in reinforcement learning, the exploration has drawn great attention in single-agent RL research. This can be even more challenging in complex environments with sparse reward [4]. In the multi-agent domain, several works [27, 19, 20] have emerged to address exploration challenges. MAVEN [27], for instance, aims to maximize the mutual information between the trajectories and latent variables, thereby enabling the learning of diverse exploration policies. EITI and EDTI [19], on the other hand, are designed to capture the influence of one agent’s behaviors on others, encouraging agents to visit states that influence the behavior of others. CMAE [20] adopts a shared exploration goal derived from multiple projected state spaces. Beyond these considerations, recent works highlight that the choice of policy class employed by the agents can significantly impact exploration dynamics [5, 6]. However, these efforts primarily concentrate on single-agent domains. In this paper, we highlight the equally crucial role of policy classes in decision-making within multi-agent systems. In our proposed approach, CPIG, we introduce an innovative policy class, the intention-guided consistency policy, specifically tailored to enhance exploration in MARL.
@@ III-B Consistency Models
The diffusion model [15] addresses the multimodal distribution matching problem using a stochastic differential equation (SDE), while the consistency model [39] tackles a comparable probability flow ordinary differential equation (PF-ODE): ${\mathrm{d}x_{\tau}}/{\mathrm{d}\tau}=-\tau\nabla\log p_{\tau}(x)$, where $\tau\in[0,T]$, $T>0$ is a fixed constant. Here, $p_{\tau}(x)=p_{\text{data}}(x)\otimes\mathcal{N}(0,\tau^{2}I)$ denotes the distribution of data $x_{\tau}$ at step $\tau$, where $\otimes$ denotes the convolution operation and $p_{\text{data}}(x)=p_{0}(x)$ represents the raw data distribution. The reverse process occurs along the solution trajectory $\{\hat{x}_{\tau}\}_{\tau\in[\epsilon,T]}$, with $\epsilon$ being a small constant close to 00, employed for handling the numerical issues at the boundary. To accelerate the sampling process of a diffusion model, the consistency model reduces the required number of sampling steps to a significantly smaller value without substantially compromising model generation performance. Specifically, it approximates a parameterized consistency function $f_{\theta}:(x_{\tau},\tau)\to x_{\epsilon}$, which is defined as a map from the noisy sample $x_{\tau}$ at step $\tau$ back to the original sample $x_{\epsilon}$, in contrast to the step-by-step denoising function $p_{\theta}(x_{\tau-1}|x_{\tau})$ as the reverse diffusion process in diffusion model. The consistency function possesses the property of self-consistency, meaning its outputs are consistent for arbitrary pairs of $(x_{\tau},\tau)$ that belong to the same PF-ODE trajectory. Besides, for modeling the conditional distribution with condition variable $c$, the consistency function is modified to $f_{\theta}(c,x_{\tau},\tau)$, representing a slight deviation from the original consistency model.
@@ IV Method
Figure 2: The overall framework of the proposed CPIG. Intention-guided diffusion policies facilitate cooperative exploration among multiple agents. The consistency policy utilizes the agent’s observation and Gaussian noise as inputs, generating actions under the guidance of masked intention. During the training phase, in addition to policy learning, the self-reference mechanism provides gradient back-propagation based on the disparity between the agent’s action and those from the reference buffer, thereby imposing a policy constraint.
In this section, we introduce CPIG, a novel approach that integrates the consistency policy guided by the shared intention between agents for exploration. As illustrated in Fig. 2, our framework consists of three principal components: a) a consistency policy that generates actions from Gaussian noise, b) a discrete intention codebook with masks, which serves as a guidance for cooperation, and c) a self-reference mechanism to constrain consistency policies. Detailed discussions on each component are provided in the subsequent sections.
@@ IV-A Consistency Policy
Notation. In this work, we distinguish between two types of timesteps, those associated with the diffusion process and those related to reinforcement learning. To clarify this distinction, we use superscripts $\tau\in[0,T]$ to represent diffusion timesteps and subscripts $t\in\{1,\dots,T_{e}\}$ to denote environment timesteps.
@@
where $\mu(\cdot,\cdot)$ representing the drift coefficient, $\sigma(\cdot)$ representing the diffusion coefficient, and $\{w^{\tau}\}_{\tau\in[0,T]}$ denoting the standard Brown Motion. Beginning with $u_{a}^{T}$, the diffusion policy aims to recover the original action $u_{a}^{0}=u_{a}$ by solving a reverse process from $T$ to 00 using the reverse-time SDE:
@@
Although this diffusion policy can be used for exploration, it is time-intensive and makes the policy inference unacceptably slow, especially in tasks that contain multiple agents. To address this issue, we introduce a faster policy, consistency policy [39, 37], based on Eq. 3:
@@
where the initial action $u_{a}^{\tau_{n}}\sim\mathcal{N}(0,{\tau_{n}}I)$, and $\theta_{a}$ represents the parameters of the policy $\pi$ of the agent $a$. We discretize the diffusion horizon $[\epsilon,T]$ into a predetermined sequence $\{\tau_{n}|n\in N\}$ of length $N$ for determining the solution trajectory of action. Moreover, $\tau_{n}$ is the $n$-th sub-intervals of the time horizon, with $n\sim\mathcal{U}(1,N-1)$. The consistency function $F_{\theta_{a}}(o_{a},u_{a}^{\tau},\tau)$ is a trainable deep neural network that takes the observation $o_{a}$ of agent $a$ as input and outputs an action with the same dimensionality as $u_{a}^{\tau}$. $c_{skip}$ and $c_{out}$ are differentiable functions, with $c_{\mathrm{skip}}(\epsilon)=1$, $c_{\mathrm{out}}(\epsilon)=0$, ensuring the consistency process remains differentiable at a fixed, small positive timestep $\epsilon$.
@@
Policy Learning. To train the consistency policy, we adopt an off-policy optimization approach similar to MADDPG [8]. Following the approach of Wang et al. [16], we integrate the Q-value function into the consistency model, training it to preferentially sample actions with higher values:
@@
IV-B Intention Guidance
To guide cooperation in multi-agent exploration, we introduce an intention learner. The intention represents an agent’s estimation of the global state, which influences its next action. By sharing the same intention, multiple agents can exhibit coordinated behaviors. We first formulate a set of $K$ discrete intentions $\Psi\coloneqq\{\psi_{1},\dots,\psi_{K}\}$, where $K\in\mathbb{N}$ is a tunable hyperparameter. Each intention $\psi_{k}$ is defined by a tuple $\left\langle e_{\psi_{k}},I_{\psi_{k}}\right\rangle$, where $k\in\{1,\dots,K\}$ is the identity of intention and $e_{\psi_{k}}\in\mathbb{R}^{m}$ is the code (or latent embedding) of intention $\psi_{k}$. $I_{\psi_{k}}$ is the set of agents who have the same intention $\psi_{k}$, and each agent can only have one intention at each timestep: $I_{\psi_{i}}\cap I_{\psi_{j}}=\varnothing,\cup_{j}I_{\psi_{j}}=I$ for $i,j\in\{1,2,\dots,K\}$ and $i\neq j$. To represent intention, we design an intention leaner with VQ-VAE [42], as shown in Fig. 3. The motivations for this include: a) The discrete codebook in VQ-VAE can effectively extract state features with similar semantics, providing better separation than traditional continuous representation methods. b) The use of discrete intention representation helps mitigate the impact of unavoidable noise on exploration in stochastic environments. c) The discretization process ensures a distinct association, where each agent is linked to a specific intention.
The intention learner comprises four primary components: an observation encoder, an intention codebook, a state decoder, and an intention mask. While the intention can theoretically be derived from the global state $s$, the CTDE paradigm imposes restrictions, allowing each agent $a$ to access only its local observation $o_{a}$ during distributed execution. Given that $o_{a}=\Omega(s,a)$ represents a projection of the global state $s$, we design a two-phase mechanism to infer intentions based on local observations. To deduce the intention from local observations, the intention learner has two modes during the training and execution phases: During the training phase, the intention learner processes observations from all agents, obtains discrete intention representations for each agent’s observation, and uses them to reconstruct the global state; In the execution phase, it processes agent $a$’s observation to obtain the intention for guiding policy generation of action $u_{a}$. In this manner, the intention learner develops the capacity to infer intention based on local observations.